CUET Mathematics: AlgebraProbability. Free, no login required.

Q1:

2026: 18 May Shift 2

Probability

Easy

Core

An alphabet has been selected from English alphabets. What is the probability that it is a vowel given that it precedes 'M' ?

Answer options
Option 4
Correct Answer
Explanation for 2026: 18 May Shift 2 MAT question 1

Q2:

2026: 18 May Shift 1

Probability

Easy

Common

A die is thrown and a coin is tossed. The probability that the outcome will be a number greater than 4 or a head or both is

Answer options
Option 1
Correct Answer
Explanation for 2026: 18 May Shift 1 MAT question 2

Q3:

2026: 18 May Shift 1

Probability

Easy

Core

Let A and B be two independent events. Then which of following statements is not true?

Answer options
Option 4
Correct Answer
Explanation for 2026: 18 May Shift 1 MAT question 3

Q4:

2026: 18 May Shift 1

Probability

Medium

Core

A bag contains 3 white and 2 black balls, another bag contains some white and 4 black balls. One bag is chosen at random. From the selected bag, one ball is drawn. If the probability of getting a black ball from the selected bag is 8/15, then the number of white balls in the second bag is

Answer options
Option 2
Correct Answer
Explanation for 2026: 18 May Shift 1 MAT question 4

Q5:

2026: 18 May Shift 1

Probability

Easy

Core

For any two events A and B, which of the following are correct?

(A) P(A∪B)=P(A)+P(B)−P(A∩B)P(A\cup B) = P(A)+P(B)-P(A\cap B)

(B) P(A)=P(A∪B)+P(A∩B)P(A) = P(A\cup B)+P(A\cap B)

(C) P(A/B)=P(A)/P(B)P(A/B) = P(A)/P(B)

(D) P(A)=P(A∩B)+P(A∩Bˉ)P(A) = P(A\cap B)+P(A\cap \bar{B})

Choose the correct answer from the options given below:

Answer options
Option 2
Correct Answer
Explanation for 2026: 18 May Shift 1 MAT question 5

Q6:

2026: 18 May Shift 1

Probability

Medium

Core

If A and B are two events such that P(A)≠0P(A) \neq 0, P(B)≠1P(B) \neq 1 and P(Aˉ/Bˉ)=1−kP(Bˉ)P(\bar{A}/\bar{B}) = \frac{1-k}{P(\bar{B})}, then the value of kk is

Answer options
Option 3
Correct Answer
Explanation for 2026: 18 May Shift 1 MAT question 6

Q7:

2026: 18 May Shift 1

Probability

Easy

Applied

The mean of the number of heads in a simultaneous toss of three coins is

Answer options
Option 4
Correct Answer
Explanation for 2026: 18 May Shift 1 MAT question 7

Q8:

2026: 15 May Shift 1

Probability

Medium

Common

Three balls are drawn one by one without replacement from a bag containing 5 white and 4 red balls. Let X denote the number of white balls.

Match the LIST-I with LIST-II

LIST-ILIST-II
A. P(X=0)P(X=0)I. 10/2110/21
B. P(X=1)P(X=1)II. 5/425/42
C. P(X=2)P(X=2)III. 1/211/21
D. P(X=3)P(X=3)IV. 5/145/14

Choose the correct answer from the options given below:

Answer options
Option 3
Correct Answer
Explanation for 2026: 15 May Shift 1 MAT question 8

Q9:

2026: 15 May Shift 1

Probability

Medium

Core

A company manufactured fans with three units A, B and C. Unit A, B and C produces 2%, 10% and 14% defective fans respectively. Units A, B and C produces 40%, 40% and 20% fans of total products respectively. A fan is randomly chosen that is found to be defective then the probability that the chosen fan is produced from unit C, is

Answer options
Option 1
Correct Answer
Explanation for 2026: 15 May Shift 1 MAT question 9

Q10:

2026: 15 May Shift 1

Probability

Medium

Core

Let A and B are two events associated to a random experiment, then

Match the LIST-I with LIST-II

LIST-ILIST-II
A. P(A∩B)+P(A∩B‾)P(A \cap B) + P(A \cap \overline{B})I. P(A∪B)P(A \cup B)
B. P(A∩B)+P(A‾∩B)P(A \cap B) + P(\overline{A} \cap B)II. P(A)+P(B)P(A) + P(B)
C. P(A∩B)+P(A‾∩B)+P(A∩B‾)P(A \cap B) + P(\overline{A} \cap B) + P(A \cap \overline{B})III. P(A)P(A)
D. P(A∪B)+P(A∩B)P(A \cup B) + P(A \cap B)IV. P(B)P(B)

Choose the correct answer from the options given below:

Answer options

Q11:

2026: 15 May Shift 1

Probability

Medium

Core

If E and F are two independent events such that P(E)=13,P(F)=35P(E)=\dfrac{1}{3}, P(F)=\dfrac{3}{5}, then which of the following is/are correct ?

A. P(E∩F)=15P(E \cap F) = \dfrac{1}{5}

B. P(E‾F)=23P\left(\dfrac{\overline{E}}{F}\right) = \dfrac{2}{3}

C. P(EF‾)=23P\left(\dfrac{E}{\overline{F}}\right) = \dfrac{2}{3}

D. P(EF)=23P\left(\dfrac{E}{F}\right) = \dfrac{2}{3}

Choose the correct answer from the options given below:

Answer options

Q12:

2026: 15 May Shift 1

Probability

Medium

Core

The probability of simultaneous occurrence of at least one of two events E and F is pp. If the probability that exactly one of E, F occurs is qq, then P(E‾)+P(F‾)P(\overline{E}) + P(\overline{F}) is equal to

Answer options

Q13:

2026: 14 May Shift 2

Probability

Easy

Common

A pair of coins thrown then

Match the LIST-I with LIST-II

LIST-ILIST-II
A.Probability of occurance of head on both coins isI.34\dfrac{3}{4}
B.Probability of occurance of at least one tail isII.12\dfrac{1}{2}
C.Probability of occurance of one tail only isIII.11
D.Probability of a tail appears on one coin given that only one coin shows head.IV.14\dfrac{1}{4}

Choose the correct answer from the options given below:

Answer options

Q14:

2026: 14 May Shift 2

Probability

Easy

Core

Four cards are drawn successively without replacement from a deck of 52 playing cards. The probability that all the four cards are king is:

Answer options

Q15:

2026: 14 May Shift 2

Probability

Medium

Core

A bag contains 5 white, 4 black and 3 red balls. Which of the following statements are True?

A. The probability of drawing three red balls one by one without replacement is 1220\frac{1}{220}

B. One by one three balls drawn without replacement then probability that third ball is red is 14\frac{1}{4}

C. The probability of drawing 2 white balls one by one without replacement is 34\frac{3}{4}

D. One by one 3 balls are drawn without replacement then probability that third ball is red is 34\frac{3}{4}

Choose the correct answer from the options given below:

Answer options

Q16:

2026: 14 May Shift 2

Probability

Easy

Core

The probability of drawing a one-hundred rupee currency from two boxes one of which contain 3 fifty-rupee currency and 2 one-hundred rupee and other box contain 2 fifty-rupee currency and 3 one-hundred rupee currency notes is-

Answer options

Q17:

2026: 14 May Shift 2

Probability

Easy

Core

A pair of unbaised dice is thrown together and sum of the numbers appearing is observed. The probability that the sum of numbers on the dice was 10, given that the observed sum was at least 9, is:

Answer options

Q18:

2026: 12 May Shift 2

Probability

Easy

Common

A die is thrown once then,

Match List-I with List-II

List-IList-II
(A) Probability of getting odd number(I) 13\dfrac{1}{3}
(B) Probability of getting number more than 4(II) 23\dfrac{2}{3}
(C) Probability of getting number less than or equal to 4(III) 12\dfrac{1}{2}
(D) Probability of getting number 2(IV) 16\dfrac{1}{6}

Choose the correct answer from the options given below:

Answer options

Q19:

2026: 12 May Shift 2

Probability

Hard

Core

A bag contains 4 balls. Let E be an event in which two balls are drawn at random and found to be white, then

Match List-I with List-II

List-IList-II
(A) Probability that all balls are white(I) 16\dfrac{1}{6}
(B) Probability of drawing 2 white ball at random given that there are 3 white ball(II) 35\dfrac{3}{5}
(C) P(E)(III) 12\dfrac{1}{2}
(D) Probability of drawing 2 white ball at random given that there is only 2 white ball(IV) 59\dfrac{5}{9}

Choose the correct answer from the options given below:

Answer options

Q20:

2026: 12 May Shift 2

Probability

Easy

Core

From a lot of 10 items containing 3 defectives, a sample of 4 items is drawn at random. Let X denotes the number of defective items in the sample. If the sample is drawn randomly, then Match List-I with List-II (P(X) denotes the probability of occurence of X)

List-IList-II
(A) P (X = 0)(I) 3C3×7C110C4\dfrac{^3C_3\times{}^7C_1}{^{10}C_4}
(B) P (X = 1)(II) 7C4×3C010C4\dfrac{^7C_4\times{}^3C_0}{^{10}C_4}
(C) P (X = 2)(III) 3C1×7C310C4\dfrac{^3C_1\times{}^7C_3}{^{10}C_4}
(D) P (X = 3)(IV) 3C2×7C210C4\dfrac{^3C_2\times{}^7C_2}{^{10}C_4}

Choose the correct answer from the options given below:

Answer options

Q21:

2026: 12 May Shift 2

Probability

Easy

Core

Suppose that A and B are independent events with P(A)=17P(A)=\dfrac{1}{7} and P(B)=37P(B)=\dfrac{3}{7}, then P(A∩B)P(A\cap B) is

Answer options

Q22:

2026: 12 May Shift 2

Probability

Easy

Core

Given that the events A and B are mutually exclusive such that P(A)=12P(A)=\dfrac{1}{2}, P(A∪B)=23P(A\cup B)=\dfrac{2}{3} and P(B)=pP(B)=p, then the value of pp is:

Answer options

Q23:

2025: 3 June Shift 2

Probability

Medium

Common

Let X denotes the number of heads in a simultaneous toss of three coins, then P(0<X<3)P(0 < X < 3) is

Answer options

Q24:

2025: 3 June Shift 2

Probability

Medium

Core

Three students A, B and C can respectively solve 50%, 25% and 20% of the problems in a book. A particular problem is selected at random from the book. The probability that at least one of them will solve the problem is

Answer options

Q25:

2025: 3 June Shift 2

Probability

Medium

Core

The probabilities of occurrence of two events E and F are 0.25 and 0.50 respectively. The probability of their simultaneous occurrence is 0.14. The probability that neither E nor F occurs is

Answer options

Q26:

2025: 3 June Shift 2

Probability

Medium

Core

A bag contains 4 red and 6 black balls. Two balls are drawn in succession without replacement. The probability that the first is red and the second is black is

Answer options

Q27:

2025: 3 June Shift 2

Probability

Medium

Core

The probability distribution of a random variable X is

X01234
P(X)0.2kk2kk

Match List-I with List-II

List-IList-II
(A) value of k(I) 1625\frac{16}{25}
(B) P(x≥2)P(x \geq 2)(II) 925\frac{9}{25}
(C) P(x=3)P(x = 3)(III) 425\frac{4}{25}
(D) P(x<2)P(x < 2)(IV) 825\frac{8}{25}

Choose the correct answer from the options given below:

Answer options

Q28:

2025: 3 June Shift 2

Probability

Medium

Applied

If X is a random variable and aa, bb are real numbers, then which of the following statements are true?

(A) Var(aX+b)=a2Var(X)Var(aX+b) = a^2 Var(X)

(B) E(aX+b)=aE(X)+bE(aX+b)= a E(X) + b

(C) E(aX+b)=aE(X)−E(b)E(aX+b)= a E(X) - E(b)

(D) Var(aX+b)=aVar(X)+bVar(aX+b)= a Var(X) + b

Choose the correct answer from the options given below:

Answer options

Q29:

2025: 3 June Shift 2

Probability

Medium

Applied

A random variable X has the following probability distribution:

X012
P(X)1/41/21/4

then, which of the following is correct?

Answer options

Q30:

2025: 3 June Shift 2

Probability

Medium

Applied

The binomial distribution for which the mean is 5 and variance 4, is

Answer options